### v3

@@ Abstract
With the growing demand for seamless connectivity and intelligent communication, the integration of artificial intelligence (AI) and sixth-generation (6G) communication networks has emerged as a transformative paradigm. By embedding AI capabilities across various network layers, this integration enables optimized resource allocation, improved efficiency, and enhanced system robust performance, particularly in intricate and dynamic environments. This paper presents a comprehensive overview of AI and communication for 6G networks, with a focus on emphasizing their foundational principles, inherent challenges, and future research opportunities. We first review the integration of AI and communications in the context of 6G, exploring the driving factors behind incorporating AI into wireless communications, as well as the vision for the convergence of AI and 6G. The discourse then transitions to a detailed exposition of the envisioned integration of AI within 6G networks, delineated across three progressive developmental stages. The first stage, AI for Network, focuses on employing AI to augment network performance, optimize efficiency, and enhance user service experiences. The second stage, Network for AI, highlights the role of the network in facilitating and buttressing AI operations and presents key enabling technologies, such as digital twins for AI and semantic communication. The final stage, AI as a Service, it is anticipated that future 6G networks will innately provide AI functions as services, supporting application scenarios like immersive communication and intelligent industrial robots. Specifically, we have defined the quality of AI service, which refers to a framework for measuring AI services within the network. We further summarize the standardization process of AI for wireless networks, highlighting key milestones and ongoing efforts. In addition, we conduct an in-depth analysis of the critical challenges faced by the integration of AI and communications in 6G. Finally, we outline promising future research opportunities that are expected to drive the development and refinement of AI and 6G communications, paving the way for the next-generation communication infrastructure.
@@
6G, AI, AI for network, AI as a service, large-scale AI models, network for AI, overview
@@ 1 Introduction
AI technology has advanced rapidly over the past decade, particularly in machine learning (ML) [4], deep learning (DL) [5], and natural language processing (NLP) [6]. These advancements have enabled AI to play a crucial role across various industries. AI, by virtue of its powerful data analysis and learning capabilities, can excavate massive data in wireless communication networks, realizing intelligent management and optimization of the network. In the 5G era, AI has been successfully applied in aspects of wireless networks, such as network optimization, traffic prediction, and fault detection, significantly enhancing network performance and user experiences [7]. However, to truly achieve AI-native support in 6G networks, many unresolved issues need to be addressed.
The introduction of AI can enhance coding efficiency by utilizing compressed semantic information (SI) to transmit more data with less bandwidth, which helps alleviate network congestion and improve data transmission rate [8]. However, AI algorithms generate additional data that needs to be transmitted, such as model parameters, training data, and real-time feedback. This raises the question of whether the overall data volume in future networks will decrease or increase. This change in data volume will directly impact network design and architecture. Moreover, both base stations (BS) and user devices will employ AI algorithms for resource allocation to reduce energy consumption and improve resource utilization. However, the operation of AI itself may increase power consumption, leading to another question: Will energy consumption in future networks increase or decrease? In the pursuit of high efficiency, it is crucial to consider optimizing energy efficiency in AI applications to ensure the sustainable development of networks. Additionally, while networks can leverage AI algorithms to enhance transmission reliability and better respond to changing network conditions and user demands, the inherent uncertainty of AI algorithms—especially in complex and dynamic network environments—raises concerns about whether future networks will operate more reliably. These uncertainties may lead to decision-making errors, affecting user experience and overall network performance.
@@
Figure 1: Organization of this article.
@@ 1.1.1 Better Utilization of Spectrum Resources
The rapid development of information technology and networks has led to a new round of business innovation and upgrading of service demand and the volume of network traffic data has reached unprecedented levels [9]. The networks face the dual challenges of explosive growth in data traffic and the demand for massive devices to connect everything. As an essential mobile communication resource, spectrum remains the most critical element in improving network capacity [10]. Whether the limited spectrum resources can be better managed is the key to whether the network can guarantee high-quality services for users. However, existing wireless communication systems still adopt a static spectrum management model, where authorized users have exclusive access. This model lacks the ability of dynamic adjustment and is unable to flexibly allocate resources like spectrum and power in accordance with the real-time requirements of services. As a result, it is liable to lead to resource waste or insufficient resources for some services, which will impact network performance and user experiences.
@@ 1.1.2 Lower-Carbon Wireless Coverage
With the widespread adoption and continuous development of 5G networks, issues related to energy consumption and costs are becoming increasingly prominent. The energy consumption of the 5G system is about three time as that of the fourth-generation (4G) system due to its wider bandwidth, more channels, more complex equipment architecture [11]. Additionally, due to the use of higher frequency bands in 5G technology, there is a further reduction in the coverage area per BS. To achieve the same coverage targets, the number of 5G BS required would be three to four times that of 4G, significantly increasing the cost of network deployment for 5G BS. Moreover, the current operational management systems of actual networks lack flexibility and intelligence. The operational status of BS cannot be dynamically adjusted in precise accordance with real-time changes in regional user traffic. During periods and in areas with low traffic demand, the BSs may maintain a high workload, leading to energy waste. The mobile communication industry in the forthcoming era is inevitably trending towards a more environmentally friendly and sustainable pathway, with continuous endeavors in the innovation of green energy-efficient technologies.
@@ 1.1.3 More Efficient and Cost-effective Network O&M
Currently, the operation and maintenance (O&M) management of mobile communication networks primarily rely on manual on-site troubleshooting, resulting in low levels of automation and high maintenance costs. This approach imposes high demands on the workload and expertise of maintenance personnel. The work mode is predominantly reactive, addressing issues as they arise rather than proactively preventing them. Consequently, providing rapid response reports and swift emergency handling capabilities is challenging. At the same time, 5G networks introduce new technologies, such as virtualization and mobile edge computing (MEC) [12]. Edge sites are rapidly expanding, characterized by their large number, wide distribution, and heterogeneity. As the number of edge sites increases, maintenance teams must expand accordingly, leading to uncontrolled labor costs [13].
Compared to conventional core network (CN) equipment’s typical millisecond-level fault detection efficiency, network resource virtualization introduces new potential fault points. Moreover, it is difficult for operators to delineate the responsibility boundaries among suppliers when faults occur. Using software mechanisms for fault detection, the virtual network function at the edge site typically experiences longer response times, which complicates timely network maintenance and management. Moreover, the incessantly emerging novel applications and services have presented elevated requisites with respect to network bandwidth, latency, reliability, and security. The conventional methodologies for network O&M management will encounter difficulties in buttressing the more intricate and mutable network milieu in the perspective. Consequently, wireless networks are obliged to probe into innovative intelligent approaches for network operation, maintenance, and management, so as to guarantee the high efficacy and promptness of network deployment, operation, and maintenance.
@@ 1.1.4 More Personalized and Customized On-demand Service Capability
In the post-5G era, the mobile communication network, as a crucial infrastructure for digital social transformation, is no longer confined to the business domain of conventional mobile communication. Instead, it places greater emphasis on fulfilling the vastly diverse new demands for digital transformation across a multitude of industries. Specifically, the application scenarios of mobile communication networks are rich and diverse, covering multiple fields such as mobile communication, mobile internet, internet of things (IoT), smart cities, and satellite internet. Different application scenarios have significant differences in requirements for network user experience, data rate, service latency, reliability, etc. The network needs to possess the ability to provide personalized and customized services to meet the user needs of different scenarios [14]. Therefore, mobile communication networks need to develop in a more intelligent and efficient direction, continuously enhancing the personalized and customized on-demand service capabilities to meet new scenarios, new services, and new demands, improving user experience, and optimizing resource utilization.
@@ 1.1.5 More Secure and Reliable Transmissions
From 2G to 5G, the design goal of mobile communication networks in terms of security and reliability has been to ensure the authenticity of user and network identities, prevent data interception and tampering during transmission, and primarily employ security mechanisms, such as network authentication, data encryption, and integrity checks, to safeguard communications. Moreover, these security measures have undergone several “add-on” enhancements and refinements. Establishing security measures is “post-hoc”, meaning that security mechanisms are added to the system after the design of the network communication functions is complete. Furthermore, some security vulnerabilities are only addressed in subsequent generations of mobile communication networks—for example, 5G remedied the capture attacks on user identity tags in 4G [15, 16]. The defense systems are deployed primarily based on a perimeter-centric defense architecture, most often erected on the premise of “known risks”. Protective devices, such as firewalls and intrusion detection systems, are positioned at the physical boundaries of mobile communication networks. Internally, these systems adopt a trusting stance, while externally, they use a “patching” approach to ward off security risks.
@@ 1.2.1 Grand Vision
As new-generation information and communication technologies and applications such as big data, cloud computing, the IoT, and AI develop, and as these intertwine deeply with information, communication, data and technology, the technology landscape transitions into an era of “digital twin (DT) and ubiquitous intelligence” in the 6G era [18]. 6G networks will build a closely interconnected network space by providing communication services that integrate the physical and virtual worlds, enabling seamless interaction between the human society, physical world, and virtual world. This integration will create new value through the digital world, realizing the promising vision of “6G changing the world.”
@@ 1.2.2 Usage Scenarios of 6G
In June 2023, the international telecommunication union (ITU) radio sector (ITU-R) approved the “Framework and Overall Objectives Recommendation for the Development of international mobile communications (IMT) for 2030 and Beyond” (referred to as “Recommendation”) at the 44th ITU-R WP 5D meeting. This document outlines the development goals, typical scenarios, and capability indicators for 6G, providing essential guidance for subsequent 6G technology and standards research. IMT-2030 (6G) has defined six significant scenarios. As shown in Figure 2, building on the “iron triangle” of IMT-2020 (5G), IMT-2030 (6G) extends outward to form a hexagon. On the outermost circle of the hexagon, four design principles applicable to all scenarios are listed: sustainability, ubiquitous intelligence, security/privacy/resilience, and connecting the unconnected.
@@
Figure 2: Usage scenarios and new capabilities of 6G proposed by IMT-2030
@@
Immersive Communication: The immersive communication scenario extends the eMBB of 5G. It includes use cases that provide users with rich interactive video (immersive) experiences, including interactions with machine interfaces. Typical use cases in this scenario include immersive extensive reality (XR) communication, remote multi-sensory presentation, and holographic communication. In immersive communication, supporting time-synchronized mixed traffic of video, audio, and other environmental data is essential, including independent support for voice. Moreover, the ability to improve spectrum efficiency, provide a consistent service experience, and balance higher data rates and enhanced mobility in various environments is crucial. Some immersive communication use cases may also require high reliability and low latency support to enable responsive and accurate interactions with real and virtual objects and greater system capacity to connect numerous devices simultaneously.
@@
Hyper-Reliable and Low-Latency Communication: The hyper-reliable and low-latency communication scenario extends the uRLLC capabilities of 5G, covering specialized use cases that are expected to have more stringent requirements for reliability and latency. Typical use cases include communications for comprehensive automation, control, and operations in industrial environments, such as robotic interactions, emergency services, remote medical care, and power transmission and distribution monitoring.
@@
Massive Communication: The massive communication scenario expands upon the mMTC capabilities of 5G, focusing on connecting many devices or sensors. Typical use cases include smart cities, transportation, logistics, healthcare, energy, environmental monitoring, agriculture, and many other fields with expanded and new applications. This scenario involves various IoT devices that may not have batteries or have long-life batteries. This scenario requires support for high connection density. Depending on the use case, the scenario requires different data rates, low power consumption, mobility, extended coverage range, and high security and reliability.
@@
Ubiquitous Connectivity: The ubiquitous connectivity scenario aims to enhance connectivity to bridge the digital division. One of the critical focuses of this use case is to address areas that currently lack coverage or have minimal coverage, especially in rural, remote, and sparsely populated areas [19, 20]. Typical use cases include, but are not limited to, IoT and mobile broadband communication in these underserved regions.
@@
AI and Communication: The AI and communication integration scenario will support distributed computing and AI-driven applications. It will enable unprecedented use cases by leveraging data collection, local or distributed computation offloading, and distributed training and inference of AI models across intelligent nodes. Typical use cases in this scenario include assisted autonomous driving, autonomous collaboration between medical devices, offloading intensive computations across devices and networks, creating and predicting DT, and assisting collaborative robots. This scenario requires support for high regional traffic capacity, user experience data rates, low latency, and high reliability tailored to specific use cases. In addition to communication aspects, this usage scenario is expected to include a range of new functionalities integrating AI and computing capabilities into IMT-2030. These include data collection, preparation, and processing from diverse sources, distributed AI model training, model sharing, and distributed inference across IMT systems, as well as coordination and linking of computing resources.
@@
Integrated Sensing and Communication: This scenario contributes to new applications and services that require sensing capabilities. It leverages IMT-2030 to provide wide-area multidimensional sensing, offering spatial information about unconnected devices and spatial information about connected devices, movement, and the surrounding environment. Typical use cases in this scenario include IMT-2030 assisted navigation, activity detection, and motion tracking (e.g., posture/gesture recognition, fall detection, vehicle/pedestrian detection), environmental monitoring (e.g., rainwater/pollution detection), and providing sensory data/information about the surrounding environment for AI, XR, and DT applications. Besides the provided communication functions, this scenario also requires support for high-precision positioning and related sensing capabilities, including distance/velocity/angle estimation, object and presence detection, localization, imaging, and mapping.

### v4

@@ Abstract
With the growing demand for seamless connectivity and intelligent communication, the integration of artificial intelligence (AI) and sixth-generation (6G) communication networks has emerged as a transformative paradigm. By embedding AI capabilities across various network layers, this integration enables optimized resource allocation, improved efficiency, and enhanced system robust performance. This paper presents a comprehensive overview of AI and communication for 6G networks, with a focus on their foundational principles, inherent challenges, and future research opportunities. We first review the integration of AI and communications in the context of 6G, exploring the driving factors behind incorporating AI into wireless communications, as well as the vision for the convergence of AI and 6G. The discourse then transitions to a detailed exposition of the envisioned integration of AI within 6G networks, divided into three progressive stages. The first stage, AI for Network, focuses on employing AI to augment network performance, optimize efficiency, and enhance user service experiences. The second stage, Network for AI, highlights the role of the network in facilitating and buttressing AI operations and presents key enabling technologies. We compare wireless network large models with conventional large language models (LLMs), and identify key design principles and components for building wireless network architectures. In the final stage, AI as a Service, it is anticipated that future 6G networks will innately provide AI functions as services, supporting application scenarios like immersive communication and intelligent industrial robots. Specifically, we define the quality of AI service, which refers to a framework for measuring AI services within the network. We further summarize the standardization process of AI for wireless networks, highlighting key milestones and ongoing efforts. In addition, we analyze the critical challenges faced by the integration of AI and communications in 6G. Finally, we outline promising future research opportunities that are expected to drive the development and refinement of AI and 6G communications.
@@
6G, AI, AI and communication, AI for network, AI as a service, LLMs, network for AI
@@ 1 Introduction
AI technology has advanced rapidly over the past decade, particularly in machine learning (ML) [4], deep learning (DL) [5], and natural language processing (NLP) [6]. These advancements have enabled AI to play a crucial role across various industries. By virtue of its powerful data analysis and learning capabilities, AI can excavate massive data in wireless communication networks, realizing intelligent management and optimization of the network. In the 5G era, AI has been successfully applied to the aspects of wireless networks, e.g., network optimization, traffic prediction, and fault detection, significantly enhancing network performance and user experiences [7]. However, many unresolved issues need to be addressed to achieve AI-native support in 6G networks.
The introduction of AI can enhance coding efficiency by utilizing compressed semantic information (SI) to transmit more data with less bandwidth, which helps alleviate network congestion and improve data transmission rate [8]. However, AI algorithms generate additional data that needs to be transmitted, such as model parameters, training data, and real-time feedback. This raises the question of whether the overall data volume in future networks will decrease or increase. This change in data volume will directly impact network design and architecture. Moreover, both base stations (BS) and user devices will employ AI algorithms for resource allocation to reduce energy consumption and improve resource utilization [9]. The operation of AI itself may increase power consumption, leading to another question: Will energy consumption in future networks increase or decrease? In the pursuit of high efficiency, it is crucial to consider optimizing energy efficiency in AI applications to ensure the sustainable development of networks. Additionally, while networks can leverage AI algorithms to enhance transmission reliability and better respond to changing network conditions and user demands, the inherent uncertainty of AI algorithms—especially in complex and dynamic network environments—raises concerns about whether future networks will operate more reliably. These uncertainties may lead to decision-making errors, affecting user experience and overall network performance.
@@
Figure 1: Organization of this article
@@ 1.1.1 Better Utilization of Spectrum Resources
The volume of network traffic data has reached unprecedented levels [10]. Networks face the challenges of explosive growth in data traffic and the demand for massive devices to connect everything. As an essential mobile communication resource, spectrum remains the most critical element in improving network capacity [11]. Whether the limited spectrum resources can be better managed is the key to ensuring that the network can provide high-quality services to users. However, existing wireless communication systems still adopt a static spectrum management model, where authorized users have exclusive access. This model lacks the ability of dynamic adjustment and is unable to flexibly allocate resources like spectrum and power in accordance with the real-time requirements of services. This likely leads to resource waste or insufficient resources for some services, which impacts network performance and user experiences.
@@ 1.1.2 Lower-Carbon Wireless Coverage
The energy consumption of 5G systems is about three times that of the fourth-generation (4G) system due to its wider bandwidth, more channels, and more complex equipment architecture [12]. Moreover, due to the use of higher-frequency bands in 5G technology, there is a further reduction in the coverage area per BS. To achieve the same coverage targets, the number of 5G BS required would be three to four times that of 4G, significantly increasing the cost of network deployment for 5G BSs. The current operational management systems of actual networks lack flexibility and intelligence, as a BS cannot be dynamically adjusted in precise accordance with real-time changes in regional user traffic. During periods and in areas with low traffic demand, the BSs may maintain a high workload, leading to energy waste.
@@ 1.1.3 More Efficient and Cost-effective Network O&M
The operation and maintenance (O&M) management of current 4G/5G networks relies primarily on manual on-site troubleshooting, resulting in low levels of automation and high maintenance costs. This approach is rather reactive, addressing issues as they arise, as opposed to proactively preventing them. Providing rapid response reports and swift emergency handling capabilities is challenging. At the same time, 5G networks introduce new technologies, such as virtualization and mobile edge computing (MEC) [13, 14]. Edge sites are expanding, characterized by their large numbers, wide distribution, and heterogeneity, resulting in expansion of maintenance teams and uncontrolled labor costs [15].
Compared to conventional core network (CN) equipment’s typical millisecond-level fault detection efficiency, network resource virtualization introduces new potential fault points. It is difficult for operators to delineate the responsibility boundaries among suppliers when faults occur. Using software mechanisms for fault detection, virtual network functions at the edge site typically experience longer response times, which complicates timely network maintenance and management. Moreover, the incessantly emerging novel applications and services have presented elevated requisites with respect to network bandwidth, latency, reliability, and security.
@@ 1.1.4 More Personalized and Customized On-demand Service Capability
In the post-5G era, a mobile communication network, as a critical infrastructure for digital social transformation, is no longer confined to the business domain of conventional mobile communication. Instead, it places greater emphasis on the vastly diverse new demands for digital transformation across a multitude of industries. Specifically, the application scenarios of mobile communication networks are rich and diverse, covering fields such as mobile communication, mobile internet, internet of things (IoT), smart cities, and satellite internet. Different application scenarios have significant differences in requirements for network user experience, data rate, service latency, reliability, etc. The network needs to possess the ability to provide personalized and customized services to meet the user needs of different scenarios [16].
@@ 1.1.5 More Secure and Reliable Transmissions
From 2G to 5G, the design goal of mobile communication networks in terms of security and reliability has been to ensure the authenticity of user and network identities, prevent data interception and tampering during transmission, and primarily employ security mechanisms, such as network authentication, data encryption, and integrity checks, to safeguard communications. Moreover, these security measures have undergone several “plug-in” enhancements and refinements. Establishing security measures is “post-hoc”, meaning that security mechanisms are added to the system after the design of the network communication functions is complete. Furthermore, some security vulnerabilities are only addressed in subsequent generations of mobile communication networks—for example, 5G remedied the capture attacks on user identity tags in 4G [17, 18]. The defense systems are deployed primarily based on a perimeter-centric defense architecture, most often erected on the premise of “known risks”. Protective devices, such as firewalls and intrusion detection systems, are positioned at the physical boundaries of mobile communication networks. Internally, these systems adopt a trusting stance, while externally, they use a “patching” approach to ward off security risks.
@@ 1.2.1 Grand Vision
As new-generation information and communication technologies and applications such as big data, cloud computing, the IoT, and AI develop, and as these intertwine deeply with information, communication, data, and technology, the technology landscape transitions into an era of “digital twin (DT) and ubiquitous intelligence” in the 6G era [20]. 6G networks will build a closely interconnected network space by providing communication services that integrate the physical and virtual worlds, enabling seamless interaction between the human society, physical world, and virtual world. This integration will create new value through the digital world, realizing the promising vision of “6G changing the world.”
@@ 1.2.2 Usage Scenarios of 6G
In June 2023, the International Telecommunication Union (ITU) radio sector (ITU-R) approved the “Framework and Overall Objectives Recommendation for the Development of International Mobile Communications (IMT) for 2030 and Beyond” [21] at the 44th ITU-R WP 5D meeting. This document outlines the development goals, typical scenarios, and capability indicators for 6G, providing essential guidance for subsequent 6G technology and standards research. IMT-2030 (6G) has defined six significant scenarios. As shown in Figure 2, building on the “iron triangle” of IMT-2020 (5G), IMT-2030 (6G) extends outward to form a hexagon. On the outermost circle of the hexagon, four design principles applicable to all scenarios are listed: sustainability, ubiquitous intelligence, security/privacy/resilience, and connecting the unconnected.
@@
Figure 2: Usage scenarios and new capabilities of 6G proposed by IMT-2030 [21]
@@
Immersive Communication. The immersive communication scenario extends the eMBB of 5G. It includes use cases that provide users with rich interactive video (immersive) experiences, including interactions with machine interfaces. Typical use cases in this scenario include immersive extensive reality (XR) communication, remote multi-sensory presentation, and holographic communication. In immersive communication, supporting time-synchronized mixed traffic of video, audio, and other environmental data is essential, including independent support for voice. Moreover, the ability to improve spectrum efficiency, provide a consistent service experience, and balance higher data rates and enhanced mobility in various environments is crucial. Some immersive communication use cases may also require high reliability and low latency support to enable responsive and accurate interactions with real and virtual objects and greater system capacity to connect numerous devices simultaneously.
@@
Hyper-Reliable and Low-Latency Communication. The hyper-reliable and low-latency communication scenario extends the uRLLC capabilities of 5G, covering specialized use cases that are expected to have more stringent requirements for reliability and latency. Typical use cases include communications for comprehensive automation, control, and operations in industrial environments, such as robotic interactions, emergency services, remote medical care, and power transmission and distribution monitoring.
@@
Massive Communication. The massive communication scenario expands upon the mMTC capabilities of 5G, focusing on connecting many devices or sensors. Typical use cases include smart cities, transportation, logistics, healthcare, energy, environmental monitoring, agriculture, and many other fields with expanded and new applications. This scenario involves various IoT devices that may not have batteries or have long-life batteries. This scenario requires support for high connection density. Depending on the use case, the scenario requires different data rates, low power consumption, mobility, extended coverage range, and high security and reliability.
@@
Ubiquitous Connectivity. The ubiquitous connectivity scenario aims to enhance connectivity to bridge the digital division. One of the critical focuses of this use case is to address areas that currently lack coverage or have minimal coverage, especially in rural, remote, and sparsely populated areas [22, 23]. Typical use cases include, but are not limited to, IoT and mobile broadband communication in these underserved regions.
@@
AI and Communication. The AI and communication integration scenario will support distributed computing and AI-driven applications. It will enable unprecedented use cases by leveraging data collection, local or distributed computation offloading, and distributed training and inference of AI models across intelligent nodes. Typical use cases include assisted autonomous driving, autonomous collaboration between medical devices, offloading intensive computations across devices and networks, creating and predicting DT, and assisting collaborative robots. This scenario requires support for high regional traffic capacity, user experience data rates, low latency, and high reliability tailored to specific use cases. In addition to communication aspects, this usage scenario is expected to include a range of new functionalities integrating AI and computing capabilities into IMT-2030. These include data collection, preparation, and processing from diverse sources, distributed AI model training, model sharing, and distributed inference across IMT systems, as well as coordination and linking of computing resources.
@@
Integrated Sensing and Communication (ISAC). This scenario contributes to new applications and services that require sensing capabilities. It leverages IMT-2030 to provide wide-area multidimensional sensing, offering spatial information about unconnected devices and spatial information about connected devices, movement, and the surrounding environment. Typical use cases include IMT-2030 assisted navigation, activity detection, and motion tracking (e.g., posture/gesture recognition, fall detection, vehicle/pedestrian detection), environmental monitoring (e.g., rainwater/pollution detection), and providing sensory data/information about the surrounding environment for AI, XR, and DT applications. Besides the provided communication functions, this scenario also requires support for high-precision positioning and related sensing capabilities, including distance/velocity/angle estimation, object and presence detection, localization, imaging, and mapping.